Circular No. 5983

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1993 RO, 1993 RP, 1993 SB, 1993 SC D. J. Tholen provides precise positions of 1993 SC, measured by M.

Senay from CCD images obtained by Tholen with the 2.2-m reflector at Mauna Kea. The reduction was carried out with the help of reference-star positions determined by O. Hainaut in the PPM system.

1993 UT  α_{2000}  δ_{2000}
Nov. 19.28252  23^{h}50^{m}47.^{s}10  −0^{o}33′12′′.1
19.41519  23 50 46.82  −0 33 13.5
20.29771  23 50 44.98  −0 33 24.5
20.42730  23 50 44.71  −0 33 26.1

Computations by B. G. Marsden indicate that a circular orbit like that on IAUC 5869 (a = 34 AU) is viable but that perihelion orbits with q = 34 to 33 AU and up to a = 44 AU can also fit. Since the best fit corresponds to a = 39 AU, a plausible assumption for assuring the object’s orbital stability is that there is a 2:3 resonance with Neptune, and the following perihelion orbit maximizes the minimum distance Δ_{N} to Neptune at 14 AU:

T = 1993 Sept. 17.795 TT  ω = 4.^{o}146
e = 0.13812  Ω = 354.646 }2000.0
q = 33.8512 AU  i = 5.159
a = 39.2759 AU n^{o} = 0.0040042 P = 246.1 years

The orbits of 1993 RO, RP and SB (IAUC 5865, 5867, 5869) can be stabilized by the same mechanism, the perihelion orbits below maximizing Δ_{N} at 13–15 AU. More eccentric “Pluto-like” stable solutions also exist.

T = 1993 Sept. 14.160 TT  ω = 179.^{o}536
e = 0.19335  Ω = 170.313 }2000.0
q = 31.6854 AU  i = 3.647
a = 39.2801 AU n^{o} = 0.0040036 P = 246.2 years
T = 1993 Sept. 15.301 TT  ω = 180.^{o}640
e = 0.11355  Ω = 192.090 }2000.0
q = 34.8632 AU  i = 2.570
a = 39.3289 AU n^{o} = 0.0039961 P = 246.6 years
T = 1993 Sept. 16.790 TT  ω = 3.^{o}598
e = 0.17630  Ω = 354.745 }2000.0
q = 32.3874 AU  i = 1.893
a = 39.3193 AU n^{o} = 0.0039976 P = 246.6 years

1994 May 3 Brian G. Marsden